# Influence of catastrophes and hidden dynamical symmetries on ultrafast backscattered photoelectrons

### v1

@@ Abstract
We discuss the effect of using potentials with a Coulomb tail and different degrees of softening in the photoelectron momentum distributions (PMDs) using the recently implemented hybrid forward-boundary CQSFA (H-CQSFA). We show that introducing a softening in the Coulomb interaction influences the ridges observed in the PMDs associated with backscattered electron trajectories. In the limit of a hard-core Coulomb interaction, the re-scattering ridges close along the polarization axis, while for a soft-core potential, they are interrupted at ridge-specific angles. We analyze the momentum mapping of the different orbits leading to the ridges. For the hard-core potential, there exist two types of saddle-point solutions that coalesce at the ridge. By increasing the softening, we show that two additional solutions emerge as the result of breaking a hidden dynamical symmetry associated exclusively with the Coulomb potential. Further signatures of this symmetry breaking are encountered in subsets of momentum-space trajectories. Finally, we use scattering theory to show how the softening affects the maximal scattering angle and provide estimates that agree with our observations from the CQSFA. This implies that, in the presence of residual binding potentials in the electron’s continuum propagation, the distinction between purely kinematic and dynamic caustics becomes blurred.
@@ I Introduction
Possibly the best-known caustics in high-order ATI are the rescattering ridges, which occur when pairs of backscattered trajectories coalesce. These ridges may extend up to a high photoelectron energy, which, for the shortest trajectory pair and linearly polarized monochromatic fields, is around $10U_{p}$, where $U_{p}$ is the ponderomotive energy. The prevalent argument, backed by classical models [23, 24], the strong-field approximation (SFA) [25] and the adiabatic approximation [26], is that the rescattering ridge is primarily a kinematic caustic. One should note, however, that, for such methods, the acts of rescattering are spatially localized and well defined, as they are either constructed for short-range potentials [26], or as a Born-type expansion around field-dressed plane waves [25]. The latter assumption restricts the influence of the rescattering potential to a single point, to which the electron returns.
@@
In the present paper, we study the rescattering ridge due to backscattered electron orbits considering binding potentials with different degrees of softening using the hybrid forward-boundary CQSFA (H-CQSFA) recently developed in [43]. We perform a comprehensive study of the role of the Coulomb singularity in obtaining closed backscattered ridges and show how using a soft-core Coulomb potential affects the maximum rescattering angle leading to open ridges. We perform the initial to final momentum mapping to separate the different sets of solutions contributing to the ridges. Using elements of catastrophe theory, we find that softening the potential gives rise to two additional sets of solutions, which are associated with breaking a dynamical symmetry that exists for the Coulomb potential but not for its soft-core counterpart. This symmetry is associated with the conservation of the Laplace-Runge-Lenz (LRL) vector. We also analyze electron trajectories corresponding to each of these branches and find that, for a hard-core potential, they are approximately circular in momentum space, which are signatures of the LRL vector being conserved. In contrast, for the soft-core potential non-Coulomb dynamics alters these features and some of these orbits become elliptical. Finally, we use scattering theory to understand how the softening parameter influences the maximum scattering angle in a field-free scenario. We show that the maximum angle observed in the rescattering ridges agrees well with the predictions of the scattering model.
@@ II Background
II.1 General CQSFA expressions
@@
where $D=\text{det}\left[\partial\mathbf{p}_{s}(t)/\mathbf{p}_{s}(t^{\prime}_{s})\right]$, and $\nu_{s}$ is the Maslov phase associated with a solution $s$ as calculated by the prescription given in [45] and [46].
@@
with a parameter $\alpha$ to soften the Coulomb singularity. In this work, we vary this parameter to assess its influence on the dynamics.
@@
where $\mathbf{E}(t)=-d\mathbf{A}(t)/dt$ is the electric field, stems from the saddle-point equation (5). The matrix element contains information about the initial bound-state geometry and also occurs in the standard strong-field approximation.
@@ II.2 Model and orbit types
where $\hat{z}$ is the polarization direction of the electric field, and $U_{p}$ is the ponderomotive energy. We assume the target to be hydrogen ($I_{p}=0.5$ a.u.) and take the electron to be initially in a $1s$ state. Information about how the prefactor is set up in this case is provided in [27]. For other targets and initial states, see, for instance, our publications [49, 40, 29, 50, 41]. Instead of using the original implementation of the CQSFA in [27, 31], which solved a purely boundary problem using as initial guesses orbits with pre-determined dynamics, in this work we use the hybrid forward-boundary version of the CQSFA (H-CQSFA) discussed in [43], in which an initial forward ensemble of orbits is launched, and they are subsequently used as guesses for a boundary problem.
@@
In [32], we have shown that a subset of these trajectories compares very closely with the SFA backscattered trajectories, which are known to lead to the ridge [25]. One of the parameters used in the classification is the rescattering time. Structurally, the SFA is a Born-type expansion that will lead to, along with the ionization time, a series of times where instantaneous interactions with the core occur: the rescattering times. In the CQSFA, the only time which is given by the saddle-point equations is the ionization time. However, the lack of a pre-defined rescattering time does not mean that there are no trajectories which behave similarly to the backscattered trajectories of the SFA. Their dynamics are such that after ionization they, travel for some time under the combined influence of the electric field and the central potential, return to a close vicinity of the core, change momentum over a very short time interval and then travel away from the core. We define an analog of the rescattering time, $t_{\textrm{resc}}$, in the CQSFA to be the time at which the position of the photoelectron is closest to the core.
@@ II.3 Focal points and caustics
The trajectories defined by the saddle point equations Eq. [(5)-(7)] exhibit focal points at times $t$ such that the Jacobian $J_{s}(t)={\partial\mathbf{p}_{s}(t)}/{\partial\mathbf{p}_{s}(t^{\prime}_{s})}$ has zero determinant. The Maslov phase $\nu_{s}$, included in Eq. (1), associated with a specific semiclassical trajectory, is the asymptotic value of a time-dependent phase factor which takes integer values and changes only at focal points of the trajectory [51]. Recently, prescriptions for computing this phase have been provided for semiclassical treatments of strong field ionization [46, 45]. Additionally, since this formulation of the CQSFA considers only trajectories confined to the plane containing the polarization vector and an additional perpendicular vector chosen without loss of generality due to the cylindrical symmetry of the Hamiltonian, additional phases may have to be inserted by hand to account for focal points which exist in the full 3-dimensional ensemble of trajectories [46, 50].
@@
To gain an intuitive understanding of what exactly is happening at a focal point of a given semiclassical trajectory, it is necessary to consider the trajectories in its immediate neighborhood as well. In Fig. 1, the momentum path of a trajectory containing a focal point is shown. For a selection of times $t$ during the trajectories’ propagation, the direction of the vectors $J_{s}(t)\hat{\mathbf{u}}$ and $J_{s}(t)\hat{\mathbf{v}}$ has been illustrated at the points $\mathbf{p}(t)$ along the momentum path where $\hat{\mathbf{u}}=[1~{}0]^{\textrm{T}}$ and $\hat{\mathbf{v}}=[0~{}1]^{\textrm{T}}$, where T indicates the transpose. For the start time, the Jacobian is equal to the identity matrix so the vectors define a square area element. For all other times which do not happen to be at focal points of the trajectory, the vectors $J_{s}(t)\hat{\mathbf{u}}$ and $J_{s}(t)\hat{\mathbf{v}}$ define a parallelogram and illustrate the linear deformation of the trajectories in the immediate neighborhood of the one illustrated in Fig. 1 under the momentum mapping at time $t$. At a focal point, the dimension of the vector space spanned by the vectors $J_{s}(t)\hat{\mathbf{u}}$ and $J_{s}(t)\hat{\mathbf{v}}$ is less than $2$. This means that the collection of trajectories which come from a neighborhood of a trajectory with a focal point as shown in Fig. 1, are focused down into a 1-dimensional line or to a single point. This picture clarifies why the existence of a focal point as $t\rightarrow\infty$ leads to a caustic, since the finite probability associated with trajectories coming from a specific area in initial momentum is squeezed into a subset of the final momentum which has zero measure.
@@
Figure 1: Schematic representation of the dynamical behavior around focal points. Panel (a) contains a momentum space path (yellow solid line) for a typical backscattered trajectory in a Coulomb potential. The capital letter I indicates the initial momentum of trajectory at the time of ionization. The red and black arrows along this path are a schematic representation of the vectors which define the deformation in time of a square area element, centered around the initial momentum of the trajectory represented by the yellow line, as the trajectories are propagated. When these vectors are linearly dependent and hence the area defined by the two vectors is zero, the trajectory has a focal point. The two focal points along this example trajectory have been indicated using a green scatter point. In panel (b), the focusing of the path around a focal point has been illustrated. An ensemble of trajectories, displaced in the direction of the eigenvector associated with the zero eigenvalue of Jacobian at the time the focal point occurs, starting from the initial momentum of the trajectory displayed in panel (a), have been plotted (scale adjusted for visibility). At the time of the focal point, the ensemble of trajectories all focus at a single point in momentum space and this bunching is illustrated more clearly in the enlarged region shown within the black dashed line.
@@ III Photoelectron momentum distributions and momentum mapping
In this section, we address how specific contributions to photoelectron momentum distributions (PMDs) relate to caustics and different sheets of solutions, populating different regions in momentum space. In Fig. 2, we plot the full photoelectron momentum distributions (PMDs) computed using the H-CQSFA for three specific values of the softening parameter $\alpha$ (see Eq. (4)). A small value, $\alpha=10^{-6}$ (Fig. 2(a)), that gives a good approximation for the exact Coulomb potential, and hence we call it hard-core Coulomb potential, and the much larger values $\alpha=10^{-4}$ (Fig. 2(b)) and $10^{-2}$ (Fig. 2(c)) to obtain a soft-core Coulomb potential. Throughout, we consider ionization times within a single field cycle in order to avoid prominent above-threshold ionization (ATI) rings arising from inter-cycle interference. These rings are irrelevant to the present work and have been studied elsewhere [56, 37, 30, 31, 28, 44].
@@
Table 1: Classical excursion times and approximate maximum kinetic energies $E_{ridge}$ at the rescattering ridges for the three shortest pairs of backscattered electron orbits, calculated for the parameters in Fig. 2. According to the H-CQSFA classification based on $\Pi_{\parallel}$ and $\Pi_{\perp}$, the first and third shortest pairs are type 4 orbits, and the second shortest pair are type 3 orbits. The rescattered times in the H-CQSFA have been determined according to our previous publication [32].
Figure 2: Photoelectron momentum distributions for atomic Hydrogen ($I_{p}=0.5$ a.u.) in the presence of an $800$ nm wavelength, $1.5\times 10^{14}~{}\textrm{Wcm}^{-2}$ intensity, linearly polarized and monochromatic field, computed using the H-CQSFA, as functions of the final momentum components $p_{f\parallel}$ and $p_{f\perp}$ parallel and perpendicular to the laser-field polarization. Panels (a), (b) and (c) correspond to Coulomb softening parameters of $\alpha=1\times 10^{-6}$, $\alpha=1\times 10^{-4}$ and $\alpha=1\times 10^{-2}$, respectively, given in red on the right top corner of the figure.

### v2

@@ Abstract
We discuss the effect of using potentials with a Coulomb tail and different degrees of softening in photoelectron momentum distributions (PMDs) using the recently implemented hybrid forward-boundary CQSFA (H-CQSFA). We show that introducing a softening in the Coulomb interaction influences the ridges observed in the PMDs associated with backscattered electron trajectories. In the limit of a hard-core Coulomb interaction, the re-scattering ridges close along the polarization axis, while for a soft-core potential, they are interrupted at ridge-specific angles. We analyze the momentum mapping of the different orbits leading to the ridges. For the hard-core potential, there exist two types of saddle-point solutions that coalesce at the ridge. By increasing the softening, we show that two additional solutions emerge as the result of breaking a hidden dynamical symmetry associated exclusively with the Coulomb potential. Further signatures of this symmetry breaking are encountered in subsets of momentum-space trajectories. Finally, we use scattering theory to show how the softening affects the maximal scattering angle and provide estimates that agree with our observations from the CQSFA. This implies that, in the presence of residual binding potentials in the electron’s continuum propagation, the distinction between purely kinematic and dynamic caustics becomes blurred.
@@ I Introduction
Possibly the best-known caustics in high-order ATI are the rescattering ridges, which occur when pairs of backscattered trajectories coalesce. These ridges may extend up to a high photoelectron energy, which, for the shortest trajectory pair and linearly polarized monochromatic fields, is around $10U_{p}$, where $U_{p}$ is the ponderomotive energy [^1] The ponderomotive energy is the time-averaged kinetic energy acquired by an electron in a strong laser field. For long enough pulses or monochromatic fields and atomic units, it is given by $U_{p}=I/(4\omega^{2})$, where $I$ and $\omega$ are the driving-field intensity and frequency, respectively. The prevalent argument, backed by classical models [23, 24], the strong-field approximation (SFA) [25] and the adiabatic approximation [26], is that the rescattering ridge is primarily a kinematic caustic. One should note, however, that, for such methods, the acts of rescattering are spatially localized and well defined, as they are either constructed for short-range potentials [26], or as a Born-type expansion around field-dressed plane waves [25]. The latter assumption restricts the influence of the rescattering potential to a single point, to which the electron returns.
@@
In the present paper, we study the rescattering ridge due to backscattered electron orbits considering binding potentials with different degrees of softening using the hybrid forward-boundary CQSFA (H-CQSFA) recently developed in [43]. We perform a comprehensive study of the role of the Coulomb singularity in obtaining closed backscattered ridges and show how using a soft-core Coulomb potential affects the maximum rescattering angle leading to open ridges. We perform the initial to final momentum mapping to separate the different sets of solutions contributing to the ridges. Using elements of catastrophe theory, we find that softening the potential gives rise to two additional sets of solutions, which are associated with breaking a dynamical symmetry that exists for the Coulomb potential but not for its soft-core counterpart. This symmetry is associated with the conservation of the Laplace-Runge-Lenz (LRL) vector. We also analyze electron trajectories corresponding to each of these branches and find that, for a hard-core potential, they are approximately circular in momentum space, which is a signature of the LRL vector being conserved. In contrast, for the soft-core potential non-Coulomb dynamics alters these features and some of these orbits become elliptical. Finally, we use scattering theory to understand how the softening parameter influences the maximum scattering angle in a field-free scenario. We show that the maximum angle observed in the rescattering ridges agrees well with the predictions of the scattering model.
@@ II Background
II.1 General Coulomb quantum-orbit strong-field approximation expressions
@@
where $D=\text{det}\left[\partial\mathbf{p}_{s}(t)/\partial\mathbf{p}_{s}(t^{\prime}_{s})\right]$, and $\nu_{s}$ is the Maslov phase associated with a solution $s$ as calculated by the prescription given in [45] and [46].
@@
with a parameter $\alpha$ to soften the Coulomb singularity. In this work, we vary this parameter to assess its influence on the dynamics. Eq. (4) represents the three dimensional soft-core potential, although, in practice, the problem is solved in the polarization plane.
@@
where $\mathbf{E}(t)=-d\mathbf{A}(t)/dt$ is the electric field, stems from the saddle-point equation (5) and $\widetilde{\mathbf{p}}_{s}=\mathbf{p}_{s}+\mathbf{A}(t^{\prime}_{s})$. The matrix element contains information about the initial bound-state geometry and also occurs in the standard strong-field approximation.
@@ II.2 Model and orbit types
where $\hat{z}$ is the polarization direction of the electric field, and $U_{p}$ is the ponderomotive energy. We assume the target to be hydrogen ($I_{p}=0.5$ a.u.) and take the electron to be initially in a $1s$ state. Information about how the prefactor is set up in this case is provided in [27]. For other targets and initial states, see, for instance, our publications [40, 29, 49, 41]. Instead of using the original implementation of the CQSFA in [27, 31], which solved a purely boundary problem using as initial guesses orbits with pre-determined dynamics, in this work we use the hybrid forward-boundary version of the CQSFA (H-CQSFA) discussed in [43], in which an initial forward ensemble of orbits is launched, and they are subsequently used as guesses for a boundary problem.
@@
In [32], we have shown that a subset of these trajectories compares very closely with the SFA backscattered trajectories, which are known to lead to the ridge [25]. One of the parameters used in the classification is the rescattering time. Structurally, the SFA is a Born-type expansion that will lead to, along with the ionization time, a series of times where instantaneous interactions with the core occur: the rescattering times. In the CQSFA, the only time which is given by the saddle-point equations is the ionization time. However, the lack of a pre-defined rescattering time does not mean that there are no trajectories which behave similarly to the backscattered trajectories of the SFA. Their dynamics are such that, after ionization, they travel for some time under the combined influence of the electric field and the central potential, return to a close vicinity of the core, change momentum over a very short time interval and then travel away from the core. We define an analog of the rescattering time, $t_{\textrm{resc}}$, in the CQSFA to be the time at which the position of the photoelectron is closest to the core.
@@ II.3 Focal points and caustics
The trajectories defined by the saddle point equations Eq. [(5)-(7)] exhibit focal points at times $t$ such that the Jacobian matrix $J_{s}(t)={\partial\mathbf{p}_{s}(t)}/{\partial\mathbf{p}_{s}(t^{\prime}_{s})}$ has zero determinant. This determinant is present in Eq. (1) as a stability factor, but here, instead of using the asymptotic limit it is useful to assess its behavior along the whole trajectory. The Maslov phase $\nu_{s}$, included in Eq. (1), associated with a specific semiclassical trajectory, is the asymptotic value of a time-dependent phase factor which takes integer values and changes only at focal points of the trajectory [50]. Recently, prescriptions for computing this phase have been provided for semiclassical treatments of strong field ionization [46, 45]. Additionally, since this formulation of the CQSFA considers only trajectories confined to the plane containing the polarization vector and an additional perpendicular vector chosen without loss of generality due to the cylindrical symmetry of the Hamiltonian, additional phases may have to be inserted by hand to account for focal points which exist in the full 3-dimensional ensemble of trajectories [46, 49].
@@
In addition to the effects associated with the final momenta and the breakdown of specific asymptotic expansions, one may associate focal points to semiclassical trajectories during their propagation. To gain an intuitive understanding of what exactly is happening at a focal point of a given semiclassical trajectory, it is necessary to consider the trajectories in its immediate neighborhood as well. In Fig. 1, the momentum path of a trajectory containing a focal point is shown, to which we have associated two arbitrary basis vectors. If the dimension of the space spanned by these vectors decreases with regard to the initial spaces dimension after a linear mapping, $J_{s}(t^{\prime})$, is performed, we have encountered a focal point at $t^{\prime}$.
@@
For a selection of times $t$ during the trajectories’ propagation, the direction of the vectors $J_{s}(t)\hat{\mathbf{u}}$ and $J_{s}(t)\hat{\mathbf{v}}$ has been illustrated at the points $\mathbf{p}(t)$ along the momentum path where $\hat{\mathbf{u}}=[1~{}0]^{\textrm{T}}$ and $\hat{\mathbf{v}}=[0~{}1]^{\textrm{T}}$, where T indicates the transpose [see Fig. 1(a)]. For the start time, the Jacobian is equal to the identity matrix so the vectors define a square area element. For all other times which do not happen to be at focal points of the trajectory, the vectors $J_{s}(t)\hat{\mathbf{u}}$ and $J_{s}(t)\hat{\mathbf{v}}$ define a parallelogram and illustrate the linear deformation of the trajectories in the immediate neighborhood of the one illustrated in Fig. 1 under the momentum mapping at time $t$. At a focal point, the dimension of the vector space spanned by the vectors $J_{s}(t)\hat{\mathbf{u}}$ and $J_{s}(t)\hat{\mathbf{v}}$ is less than $2$. This means that the collection of trajectories which come from a neighborhood of a trajectory with a focal point as shown in Fig. 1(b), are focused down into a 1-dimensional line or to a single point. This picture clarifies why the existence of a focal point as $t\rightarrow\infty$ leads to a caustic, since the finite probability associated with trajectories coming from a specific area in initial momentum is squeezed into a subset of the final momentum which has zero measure. In the present problem, an example of a focal point is when the momentum component changes sign due to the influence of the binding potential. This is a focal point because neighboring trajectories will undergo a bunching once this change occurs. However, this is not the only physical mechanism leading to focal points.
@@
Figure 1: Schematic representation of the dynamical behavior around focal points. Panel (a) contains a momentum space path (yellow solid line) for a typical backscattered trajectory in a Coulomb potential. The capital letter I indicates the initial momentum of trajectory at the time of ionization. The red and black arrows along this path are a schematic representation of the vectors which define the deformation in time of a square area element, centered around the initial momentum of the trajectory represented by the yellow line, as the trajectories are propagated. When these vectors are linearly dependent and hence the area defined by the two vectors is zero, the trajectory has a focal point. The two focal points along this example trajectory have been indicated using a green scatter point. In panel (b), the focusing of the path around a focal point has been illustrated. An ensemble of trajectories, displaced in the direction of the eigenvector associated with the zero eigenvalue of the Jacobian at the time at which the first focal point occurs, starting from the initial momentum of the trajectory displayed in panel (a), have been plotted (scale adjusted for visibility). At the time of the focal point, the ensemble of trajectories all focus at a single point in momentum space and this bunching is illustrated more clearly in the enlarged region shown within the black dashed line.
@@ III Photoelectron momentum distributions and momentum mapping
In this section, we address how specific contributions to photoelectron momentum distributions (PMDs) relate to caustics and different sheets of solutions, populating different regions in momentum space.
III.1 Photoelectron momentum distributions
In Fig. 2, we plot the full photoelectron momentum distributions (PMDs) computed using the H-CQSFA for three specific values of the softening parameter $\alpha$ (see Eq. (4)). A small value, $\alpha=10^{-6}$ (Fig. 2(a)), that gives a good approximation for the exact Coulomb potential, and hence we call it hard-core Coulomb potential, and the much larger values $\alpha=10^{-4}$ (Fig. 2(b)) and $10^{-2}$ (Fig. 2(c)) to obtain a soft-core Coulomb potential. Throughout, we consider ionization times within a single field cycle in order to avoid prominent above-threshold ionization (ATI) rings arising from inter-cycle interference. These rings are irrelevant to the present work and have been studied elsewhere [55, 37, 30, 31, 28, 44].
@@ III.1 Photoelectron momentum distributions
Table 1: Classical excursion times and approximate maximum kinetic energies $E_{ridge}$ at the rescattering ridges for the three shortest pairs of backscattered electron orbits, calculated for the parameters in Fig. 2. According to the H-CQSFA classification based on $\Pi_{\parallel}$ and $\Pi_{\perp}$, the first and third shortest pairs are type 4 orbits, and the second shortest pair are type 3 orbits. In principle, these energies apply for both hard- and soft-core potentials. However, for the soft potential, these values do not hold for all angles due to the additional constraint imposed by the softening. The rescattered times in the H-CQSFA have been determined according to our previous publication [32].
Figure 2: Photoelectron momentum distributions for atomic Hydrogen ($I_{p}=0.5$ a.u.) in the presence of an $800$ nm wavelength, $1.5\times 10^{14}~{}\textrm{Wcm}^{-2}$ intensity, linearly polarized and monochromatic field, computed using the H-CQSFA, as functions of the final momentum components $p_{f\parallel}$ and $p_{f\perp}$ parallel and perpendicular to the laser-field polarization. Panels (a), (b) and (c) correspond to Coulomb softening parameters of $\alpha=1\times 10^{-6}$, $\alpha=1\times 10^{-4}$ and $\alpha=1\times 10^{-2}$, respectively, shown in the upper right corner of the figure panels.
@@
For a full TDSE computation, performed with the freely available software Qprop [56], the effect of the ridges remaining open up to a specific angle is also present although it is more subtle. This is illustrated in Fig. 4, for which one can see a backscattered ridge for small softening [Fig. 4(a)], and a strong suppression for the outermost ridge if a larger softening parameter is taken [Fig. 4(b)].
The difference in subtlety is possibly due to several causes. First, the constraints associated with the ridges are more blurred, so that they are more visible in the CQSFA. Second, the initial wave packet has a width that will influence the electron-momentum distributions and the patterns observed: narrower wave packets will favor scattering ridges and broader wave packets will probe holographic structures such as the spider and the fan [43]. We can predetermine this in the CQSFA, but in TDSE computations, there are other factors, such as the influence of excited bound states and bound-state depletion. Third, because the TDSE is not an orbit-based method, it is not straightforward to disentangle the ridges associated with different electron returns and also to single out other structures such as holographic patterns. Fourth, it is easier to probe the behavior near the potential minimum in the CQSFA than in the TDSE due to the orbits providing a precise pathway, while the returning wavepacket will exhibit a larger degree of uncertainty.
Figure 4: Photoelectron momentum distributions calculated with the freely available software Qprop using the same field parameters as in Fig. 2 and softening parameters $\alpha=10^{-6}$, $\alpha=10^{-2}$ (top and bottom panel, respectively). For clarity, the ridge for the hard-core potential is indicated by the dotted white line. The orange line is meant as guidance to facilitate seeing that the ridges change.
III.2 Momentum mapping and catastrophes